Eliminate θ from the equations 4x = cos(3θ) + 3 cos θ; 4y = 3 sin θ  sin(3θ)
1
Expert's answer
2020-05-06T19:58:39-0400

Eliminate "\theta" from the equations

"4x=cos(3\theta)+3cos\theta; 4y=3sin\theta-sin(3\theta)"

Solution:

"cos(3\theta)=4cos^3\theta-3cos\theta"

"sin(3\theta)=3in\theta-4cos^3\theta"

"4x=4cos^3\theta-3cos\theta+3cos\theta"

"4y=3sin\theta-3sin\theta+4sin^3\theta"

"x=cos^3\theta"

"y=sin^3\theta"

"x^{\frac{1}{3}}=cos\theta"

"y^{\frac{1}{3}}=sin\theta"

"cos^2\theta+sin^2\theta=1"

Aswer:

"x^{\frac{2}{3}}+y^{\frac{2}{3}}=1."